A large chunk of ice with mass of 15 kg falls from a roof 8 m above the ground. What is the impact speed of the ice when it hits the ground?


1
Expert's answer
2021-05-05T14:08:45-0400

The potential energy of the chunk of ice is transformed into its kinetic energy:

"\displaystyle mgh = \frac{mv^2}{2}"

"\displaystyle gh = \frac{v^2}{2}"

"\displaystyle v = \sqrt{2gh} = \sqrt{2 \cdot8 \cdot 9.81} = 12.528" m/s.

The impact speed of falling body doesn't depend on its mass.